Find all numbers at which f is not continous for the given functions. Answers must be written in interval notation.

a. g(x) = x+5/x²+6x+8

b. f(x) = (3/x²) + (2/x+4) - (x/2x-3)

c. f(x) = x/(x³-x²)

d. f(x) = [(sqrtx)+10]/x

e. f(x) = 2x/x(sqrt x+4)

f. f(x) = ln srqt(x+6/4x²-9)

g. f(x) = ln srqt(x-8/x+3)

h. f(x) = [(sqrtx)+5]/4-(sqrt x)

i. f(x) = x-1/(sqrt x²-3x+2)

j. f(t) = sin^-1 (2t-3)


1
Expert's answer
2021-11-08T06:19:16-0500

a. "g(x) = {\frac {x+5} {x²+6x+8}}"

D: "x \in R" / {-2, -4}

g(x) is not continuous at x є {-2, -4}


b. "g(x) = {\frac 3 {x²}}+{\frac 2 {x+4}}+{\frac x {2x-3}}"

D: "x \in R" / {0, -4, 1.5}

g(x) is not continuous at x є {0, -4, 1.5}


c. "g(x) = {\frac x {x^{3}-x^{2}}}"

D: "x \in R" / {0, 1}

g(x) is not continuous at x є {0, 1}


d. "g(x) = {\frac {sqrt(x)+10} x}"

D: "x \in (0,+\infty)"

g(x) is not continuous at x є "(-\infty,0]"


e. "g(x) = {\frac {2x} {x*sqrt(x+4)}}"

D: "x \in (-4,+\infty)" / {0}

g(x) is not continuous at x є "(-\infty,-4] ∪"{0}


f. f(x) = "ln {\frac {x+6} {4x^2-9}}"

D: "x \in (-6,-1.5)∪(1.5, +\infty)"

f(x) is not continuous at x є "(-\infty,-6] ∪[-1.5,1.5]"


g. f(x) = "ln{\frac {x-5} {x+3}}"

D: "x \in (-\infty,-3)∪(8,+\infty)"

f(x) is not continuous at x є "[-3,8]"


h. f(x) = "{\frac {sqrt(x+5)} {4-sqrt(x)}}"

D: "x \in (0,+\infty)" \ {4}

f(x) is not continuous at x є "(-\infty,0] ∪"{4}


i. f(x) = "{\frac {x-1} {sqrt(x^2-3x+2)}}"

D: "x \in R" \ [1;2]

f(x) is not continuous at x є "[1,2]"


j. f(t) = "ln{\frac 1 {sin(2t-3)}}"

D: "t \in R" \ {"{\frac {2\pi n+3} 2},n \in Z" }

f(t) is not continuous at t є {"{\frac {2\pi n+3} 2},n \in Z" }